OpenStudy (anonymous):

x2-6x+5=0 solution to the equation thanks

OpenStudy (anonymous):

(x-5)(x-1) = 0, x = 5, 1

OpenStudy (anonymous):

can i see it step by step please

OpenStudy (anonymous):

x^2 - 6x + 5 = 0 so factorise this using two numbers which add to give -6 and multiply to give +5 (x-5)(x-1) = 0

OpenStudy (anonymous):

okidoki thnaks duck